OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6) 7 -14 0 14 Sorry i forgot the answer choices.^^

OpenStudy (catlover5925):

the answer is 14...if you graph it start from the bottom point and count how every many up the over so it would be 14/1 or 14

OpenStudy (kiakat):

-14

OpenStudy (anonymous):

Why is it negative 14.? @KiaKat

OpenStudy (kiakat):

because the line goes down

OpenStudy (catlover5925):

oh yea it would be -14 because you have to start from ur first point and the line slopes down..positive slopes always go up and to the right

OpenStudy (kiakat):

like it goes from the top left towards the bottom right. if it was going from the bottom left to the top right it would be positive.

OpenStudy (anonymous):

8-(-6) = 8+6 which equals 14. then -4-(-3) = -4+3 which equals -1. now you have 13/-1. one negative with one positive equals a negative, so the answer is -14.

OpenStudy (anonymous):

Okay, thank you.